OpenStudy (okdutchman7):

Please help. WILL MEDAL! The wheel was rolling along at 30 radians per second. What was its linear velocity in miles per hour if the radius was 6 meters? @mathstudent55 @satellite73 @jim_thompson5910 @nincompoop @jigglypuff314 @sleepyjess @zepdrix @Zarkon @iPwnBunnies

OpenStudy (okdutchman7):

\[\frac{ 30 rad }{ 1 \sec }\]

OpenStudy (mathstudent55):

\(\omega = 30 ~\dfrac{rad}{sec} \) \(v = r\omega \)

OpenStudy (mathstudent55):

\(r = 6 ~m\)

OpenStudy (okdutchman7):

v = velocity? r = radius?

OpenStudy (mathstudent55):

Use the formula for linear velocity with your given info. Then convert units.

OpenStudy (mathstudent55):

Yes. \(\omega\) = angular velocity \(v\) = linear velicity \(r\) = radius

OpenStudy (mathstudent55):

^velocity

OpenStudy (mathstudent55):

Sorry, but gtg. You're in good hands with Jim writing something.

OpenStudy (okdutchman7):

\[v = 6m (\frac{ 30 rad }{ \sec })\]

jimthompson5910 (jim_thompson5910):

Start with \[\Large \frac{ 30 \text{ rad} }{ 1 \text{ sec} }\] Then multiply by the conversion factor (C meters/2pi radians) \[\Large \frac{ 30 \text{ rad} }{ 1 \text{ sec} } \times \frac{ C \text{ meters} }{ 2\pi \text{ rad} }\] where C is the circumference of the circle Notice how the radian units cancel. This leaves just the units "meters" and "seconds" \[\Large \frac{ 30 {\color{red}{\cancel{{\color{black}{\text{ rad}}}}} }}{ 1 \text{ sec} } \times \frac{ C \text{ meters} }{ 2\pi {\color{red}{\cancel{{\color{black}{\text{ rad}}}}} }}\]

jimthompson5910 (jim_thompson5910):

Then you'll use conversion factors to go from meters/second to miles/hour

OpenStudy (okdutchman7):

\[\frac{ 64800000 }{ 321868.8\pi }\frac{ mi }{ hr }\]

jimthompson5910 (jim_thompson5910):

I'm not sure how you have pi in your answer. The pi terms should cancel. I would leave C in terms of pi then allow the pi terms to cancel.